• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Carlos

    Log entry time 15:48:32 on November 1,2004

    Entry number 135563

    keyword=Day shift summary

    Shift crew:
    ============
    C. Munoz Camacho
    J. Roche
    R. Gilman

    Shift summary:
    ===============

    08:00 Shift started with no beam in the hall.

    08:45 MCC called. Beam is back. Restarting DVCS production.

    09:30 Compton cavity was fliping, but rates were zero. Called Sirish. Coda restart. Everything look ok with Compton right now.

    11:50 MCC called. Hall C will start running at 110uA around noon. Also, at 1:30 the beam will be taken away for accelerator maintenance.

    12:20 Bleedthrough measurement while Hall C running at 98uA: 4.52%. Asked MCC to try and reduce it.

    12:24 Bleethrough got down to 2.50% by MCC.

    12:30 Hall C tests at high current are finished.

    12:39 Hall C back to normal. Bleedthrough measurement: 2.14%

    12:41 TWOARM run on LH2.

    12:50 Back to DVCS production.

    01:30 Beam taken away...

    ======

    Run 4279 during Hall C high current tests.